Problem 2-17. service Company use of Predetermined overhead rate [lo 4, 5, 6] Southwest Community Hospital uses a job-order costing system to track charges for each patient. The following charges relate to Candice O’Dea, who had laparoscopic knee surgery:

Pharmacy $450

Sterile supply 1,500

Supplies, other 4,500

Operating room service (nursing staff) 4,500

Anesthesia 1,500

Anesthesiologist 3,500

Southwest charges patients for the use of the operating room based on a predetermined overhead rate. The estimated overhead costs of the operating room are $5,500,000, and the activity base is available hours of 2,000. Note that the hospital does not charge for the services of the surgeon. These charges are billed directly by the surgeon.

a. Calculate the predetermined overhead rate for the operating room, and determine the charge to Candice. Assume that her surgery took 3 hours in the operating room.

b. Calculate the total cost for the knee surgery.
Answer #1

a. Calculate the predetermined overhead rate for the operating room, and determine the charge to Candice. Assume that her surgery took 3 hours in the operating room.

Predetermined overhead rate = $5,500,000 / 2,000 = 2750/hr

Applied overhead rate = 2750 * 3= $8250

b. Calculate the total cost for the knee surgery.

Total Charge = Direct Labor (Anesthesiologist + nursing staff) + Direct Material (Pharmacy + Sterile supply + Supplies, other + Anesthesia) + MOH

= (3500 + 4500) + (450 + 1500 + 4500 +1500) + 8250 = $23,150